# Features of the EAR2 neutron beam following the spallation target upgrade at the n_TOF facility at CERN

### v1

@@ 4.3 Comparison with the previous spallation target
As a result of the redesign and installation of the new spallation target, a general increase in the absolute value of the neutron flux can be observed in the whole energy range. The results of the neutron flux evaluation for the previous target (2014-2018) [29], are compared to the one with the new target (2021-present) in Figure 9. In particular, an increase at average of 40% is observed between 1 eV and 100 keV, while of about 20% below 0.5 eV. From 100 keV to 10 MeV the increase is 50%. This further increases the performance of measurements with small mass samples and/or small cross sections, and improve the signal to background ratio for very radioactive samples at EAR2, with respect to the previous spallation target.
@@ 5 Spatial beam profile
Figure 11 then shows the neutron beam profile for two different energy ranges: from 10 meV to 100 keV, and from 100 keV up to 100 MeV. Both energy ranges show a beam profile (with approximately a Gaussian shape) with 3 cm FWHM. However, the actual beam shape at these two energy ranges differs. This is due to the strong directionality of high-energy neutrons, mostly aligned with the direction of the proton beam, whereas lower-energy neutrons, due to moderation processes, have a more homogeneous distribution.
Figure 12 shows a comparison of the projections along the and axes at the maximum of the neutron beam profile between PPACMon data and FLUKA simulations integrated over the whole energy range. The asymmetries in the projection along the axis show that the beam does not have exactly a Gaussian shape. The simulations are in good agreement with the experimental data.
@@
Indeed, the accurate determination of the spatial profile is relevant, in particular, in the cross section measurement of samples which size is smaller than the neutron beam. In this case, only a fraction of the neutron beam is intercepted. This fraction is typically called the beam interception factor (BIF). The top panel of Figure 14 shows the experimental BIF (lines with markers) from PPACMon measurement for samples of different diameter, and aligned with the neutron beam centre experimentally determined at thermal energy. The solid lines represent the BIF calculated with FLUKA for the same diameters. The larger BIF between 10 and 100 meV than for higher neutron energies is due to the choice of neutron energy range for the sample alignment. The bottom panel of Figure 14 presents the ratios between PPACMon and FLUKA, showing a sufficiently good agreement that can be exploited for the measurement planning. However, these differences must be considered if simulations are used to analyse measurements sensitive to the beam profile.
@@ 6 Energy resolution
A common feature of neutron TOF facilities is the fact that all the neutrons of a given energy do not exit the target-moderator assembly at the same time, thus making the time-to-energy relation non-univocal. As introduced in Sec. 4, the relation between these two quantities is known as the energy resolution function (RF) of the facility and it can only be determined by means of MC simulations, which are then validated with experimental data.
@@
To illustrate the impact of the RF in the broadening of the resonances, in the top panel of Figure 17 the relative intrinsic widths of the 197Au(n,) resonances, , in the JEFF-3.3 library (black dots) are compared to the experimental broadening contributions corresponding to the relative width at half maximum (FWHM/). In this figure, the overall resonance broadening is displayed as a thick (blue) solid line, while the individual components (Doppler effect and RF) are displayed as dashed lines. The Doppler broadening is the main contribution at neutron energies eV, while the resolution broadening becomes the most important at higher energies. The total width for the observed 197Au(n,) resonances (blue dots) is fully dominated by the experimental broadening for eV.
@@
To illustrate the resolving power at EAR2, the bottom panel of Figure 17 shows the energy spacing between neighbouring 197Au(n,) resonances from the JEFF-3.3 library divided by the energy of the first of these resonances (black dots). The blue curve in this figure represents the total experimental broadening, which is representative of the total resonance width for 197Au. Known resonances below the blue line are separated by an energy smaller than the observed resonance width, and thus are overlapped with their neighbour resonance. To indicate the impact of this overlap we added the orange straight line to the figure. For each neutron energy, the line corresponds to , for which 50% of all the spacings lie below the line according to the expected Wigner distribution of nearest s-wave resonance spacings (D0 = 15.5 eV [34]). The energy of the intersection between the orange lines and the blue curve, around 1.5 keV, then indicates the energy at which about a half of the resonances will remain unresolved from their neighbouring resonances. As indicated by the data of the experimental commissioning, this neutron energy really represents the upper limit for the analysis of resolved 197Au(n,) resonances in n_TOF-EAR2.
@@ 7 Conclusions
The FLUKA simulations were also used to calculate the energy resolution function of the facility. The resolution function of the EAR2 beamline is significantly narrower compared to the previous spallation target. The use of this calculated resolution function in the SAMMY R-matrix code perfectly reproduces the resonance shapes observed in the 197Au(n,) and 56Fe(n,) experimental data for neutron energies up to tens of keV. This improvement will allow to determine resonance parameters of observed resonances with a higher precision and over a wider energy range than with the previous spallation target. If we consider the reference neutron capture reaction on 197Au, the resolution function available at EAR2 should now allow resolving of about half of the resonances up to a neutron energy of about 1.5 keV.
The new features of the n_TOF-EAR2 facility presented here, along with the advanced detector systems currently under development [32, 36], provide an excellent foundation for conducting new and more challenging physics measurements. Notable examples include the first-ever cross section measurements on mg samples of radioactive isotopes such as 94Nb(n,) [37] and 79Se(n,) [38], as well as new measurements of the 146Nd(n,) [39] and 243Am(n,f) cross sections [40], among numerous other ongoing measurements included in the experimental programme of the facility.

### v2

@@ 4.3 Comparison with the previous spallation target
As a result of the redesign and installation of the new spallation target, a general increase in the absolute value of the neutron flux can be observed in the whole energy range. The results of the neutron flux evaluation for the previous target (2014-2018) [31], are compared to the one with the new target (2021-present) in Figure 9. In particular, an increase at average of 40% is observed between 1 eV and 100 keV, while of about 20% below 0.5 eV. From 100 keV to 10 MeV the increase is 50%. This further increases the performance of measurements with small mass samples and/or small cross sections, and improve the signal to background ratio for very radioactive samples at EAR2, with respect to the previous spallation target. A comparison with the neutron flux in EAR1 using the new target is planned for a forthcoming work, currently under preparation, focused on the MC simulations of both areas.
@@ 5 Spatial beam profile
Figure 11 then shows the neutron beam profile for two different energy ranges: from 10 meV to 100 keV, and from 100 keV up to 100 MeV. Both energy ranges show a beam profile (with approximately a Gaussian shape) with 3 cm FWHM. However, the actual beam shape and centre at these two energy ranges differs. This is due to the strong directionality of high-energy neutrons, mostly aligned with the direction of the proton beam, whereas lower-energy neutrons, due to moderation processes, have a more homogeneous distribution.
Figure 12 shows a comparison between PPACMon data and FLUKA simulations, of the projections along the and axes at the peak value of the neutron beam profile, integrated over the whole energy range. The asymmetries in the projection along the axis show that the beam does not have exactly a Gaussian shape. The simulations are in good agreement with the experimental data.
@@
Indeed, the accurate determination of the spatial profile is relevant, in particular, in the cross section measurement of samples which size is smaller than the neutron beam. In this case, only a fraction of the neutron beam is intercepted. This fraction is typically called the beam interception factor (BIF). The top panel of Figure 14 shows the experimental BIF (lines with markers) determined with PPACMon for samples of different diameters, aligned with the beam centre at thermal energy. The solid lines represent the BIF calculated with FLUKA for the same diameters. The larger BIF between 10 and 100 meV than for higher neutron energies is due to the choice of neutron energy range for the sample alignment. The bottom panel of Figure 14 presents the ratios between PPACMon and FLUKA, showing a sufficiently good agreement that can be exploited for the measurement planning. However, these differences must be considered if simulations are used to analyse measurements sensitive to the beam profile.
@@ 6 Energy resolution
A common characteristic of neutron TOF facilities is that neutrons of a given energy do not all exit the target-moderator assembly simultaneously, making the time-to-energy relation non-unique. As introduced in Sec. 4, the relation between these two quantities is known as the energy resolution function (RF) of the facility and it can only be determined by means of MC simulations, which are then validated with experimental data.
@@
To illustrate the impact of the RF in the broadening of the resonances, in the top panel of Figure 17 the relative intrinsic widths of the 197Au(n,) resonances, , in the JEFF-3.3 library (black dots) are compared to the experimental broadening contributions corresponding to the relative width at half maximum (FWHM/). In this figure, the overall resonance broadening is displayed as a thick (green) solid line, while the individual components (Doppler effect and RF) are displayed as dashed lines. The Doppler broadening is the main contribution at neutron energies eV, while the resolution broadening becomes the most important at higher energies. The total width for the observed 197Au(n,) resonances (green dots) is fully dominated by the experimental broadening for eV.
@@
To illustrate the resolving power at EAR2, the bottom panel of Figure 17 shows the energy spacing between neighbouring 197Au(n,) resonances from the JEFF-3.3 library divided by the energy of the first of these resonances (black dots). The green curve in this figure represents the total experimental broadening, which is representative of the total resonance width for 197Au. Known resonances below the green line are separated by an energy smaller than the observed resonance width, and thus are overlapped with their neighbour resonance. To indicate the impact of this overlap we added the orange straight line to the figure. For each neutron energy, the line corresponds to , for which 50% of all the spacings lie below the line according to the expected Wigner distribution of nearest s-wave resonance spacings (D0 = 15.5 eV [36]). The energy of the intersection between the orange lines and the green curve, around 1.5 keV, then indicates the energy at which about a half of the resonances will remain unresolved from their neighbouring resonances. As indicated by the data of the experimental commissioning, this neutron energy really represents the upper limit for the analysis of resolved 197Au(n,) resonances in n_TOF-EAR2.
@@
The newly upgraded resolution of n_TOF-EAR2 is compared to other TOF facilities worldwide featuring white neutron beams and similar flight paths in Table 4. This table summarises for all the listed facilities the relative energy resolution (FWHM) at two neutron energies, 1 eV and 1 MeV. The results show how after the upgrade, n_TOF-EAR2 presents one of the bests resolutions at low neutron energies (1 eV) among the facilities with similar flight paths. The values indicated for DANCE will be significantly improved after the recent upgrade of the spallation target to Mark-IV [38]. As for ANNRI and CNSN back-n, the values in Table 4 do not consider the broadening due to the double bunch structure.
| Facility | L (m) | =1 eV | =1 MeV |
| n_TOF-EAR2 | 20 | 2.8 | 5.8 |
| GELINA [39] | 32 | 1.3 | 3.5 |
| DANCE [38] | 20 | 3.5 | 2.5 |
| ANNRI [40] | 21.5 | 3.5 | 9.0 |
| CNSN BACK-N [41] | 56 | 4.0 | 2.2 |
| n_TOF-EAR1 [4] | 185 | 3.2 | 5.4 |
@@ 7 Conclusions
The FLUKA simulations were also used to calculate the energy resolution function of the facility. The resolution function of the EAR2 beamline is significantly narrower compared to the previous spallation target. The use of this calculated resolution function in the SAMMY R-matrix code perfectly reproduces the resonance shapes observed in the 197Au(n,) and 56Fe(n,) experimental data for neutron energies up to tens of keV. This improvement will allow to determine resonance parameters of observed resonances with a higher precision and over a wider energy range than with the previous spallation target. If we consider the reference neutron capture reaction on 197Au, the resolution function available at EAR2 should now allow resolving individual resonances up to a neutron energy of about 1.5 keV. After the recent upgrade, n_TOF-EAR2 presents one of the best resolutions at low neutron energies (e.g. 1 eV) among the TOF facilities with comparable flight paths.
The new capabilities of the n_TOF-EAR2 facility presented here, along with the advanced detector systems currently under development [34, 42], provide an excellent foundation for conducting new and more challenging physics measurements. Notable examples include the first-ever cross section measurements on mg samples of radioactive isotopes such as 94Nb(n,) [43] and 79Se(n,) [44], as well as new measurements of the 146Nd(n,) [45] and 243Am(n,f) cross sections [46], among numerous other ongoing measurements included in the experimental programme of the facility.